The equation of the line which contains the point (5,10) and is perpendicular to the line through(5,10) and (25,0) has y–intercept (0, b). What is b?
Answer by Alan3354(69443)

The equation of the line which contains the point (5,10) and is perpendicular to the line through(5,10) and (25,0) has y–intercept (0, b). What is b?
------------------
Find the slope, m, of the line thru the 2 points.
m = diffy/diffx = -10/20 = -1/2
---------------------
The slope of lines perpendicular is the negative inverse, = +2
-----------
Use y = mx + b and the point (5,10) to find b:
10 = 2*5 + b
b = 0